# Forecasts for Galaxy Formation and Dark Matter Constraints from Dwarf Galaxy Surveys

### v1

@@ 6.3 Interpretation
Our results indicate that future dwarf galaxy surveys can set upper limits on the SHMF at peak subhalo masses from $10^{9}$ to $10^{10}~{}M_{\mathrm{\odot}}$, while simultaneously placing a robust lower limit at $10^{8}~{}M_{\mathrm{\odot}}$. The upper limits on the SHMF amplitude at high masses result from the fact that, if high-mass subhalos are more abundant than in the input CDM model, observed luminosity functions are over-predicted. In particular, it is difficult to “hide” the galaxies that occupy high-mass subhalos by varying galaxy–halo connection parameters without drastically reducing the abundance of lower-mass galaxies and resulting in worse agreement with the mock CDM data (for details, see our analysis of the full SHMF forecast posteriors in Appendix C.3). At lower peak halo masses, and particularly near $M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$ where our fiducial galaxy formation cutoffs take effect, models with suppressed low-mass subhalo abundances under-predict faint-end luminosity functions; however, models with enhanced low-mass subhalo abundances can be “hidden” by the galaxy formation cutoff. More precise constraints on the mass scale and shape of the cutoff would therefore improve our projected SHMF enhancement measurements.
@@
Our SHMF forecasts can be translated to any DM or early-universe physics scenario that suppresses or enhances the SHMF relative to CDM. This is most often realized by suppressing or enhancing the formation of (sub)halos relative to CDM; it will therefore be interesting to generalize our modeling framework to directly constrain the linear matter power spectrum, which often dictates the resulting suppression or enhancement of subhalo abundances in such scenarios (e.g., Gilman et al. 2022). Meanwhile, if the SHMF is altered at late times (e.g., via subhalo disruption due to new DM physics), our forecasts are applicable if the abundance of satellite galaxies is correspondingly altered. The connection between subhalo and satellite disruption is non-trivial and requires detailed study in the context of specific DM models (e.g., see Nadler et al. 2020a, 2021a and Mau et al. 2022 for examples in the context of self-interacting and decaying DM); we leave studies of such scenarios to future work.
@@
Our SHMF suppression results are relevant for a variety of warm, interacting, and ultra-light DM models (e.g., see Bechtol et al. 2022). Meanwhile, DM models that enhance the SHMF more easily evade constraints from dwarf galaxy population statistics because their effects can be counteracted by galaxy formation physics. Such models include, e.g., ultra-light axions that undergo parametric resonance in the early universe (Arvanitaki et al., 2020; Cyncynates et al., 2022) or DM produced during a period of early matter domination (Erickcek & Sigurdson, 2011). We expect that dwarf galaxy observables beyond luminosity, size, and distance (e.g., velocity dispersions; Kim & Peter 2021; Esteban et al. 2023) will help probe DM physics that enhances (sub)halo abundances. Distinct small-scale structure probes like strong gravitational lensing and stellar streams, which are more sensitive to subhalos density profiles than dwarf galaxy population statistics, will complement this effort.
@@ 7.1 Near-Term Observational Capabilities
Our assumption that complete satellite populations of several MW-mass systems can be detected is clearly optimistic given current and near-future observational capabilities. In particular, this level of completeness will be difficult to achieve even for the MW and M31 in the next decade. Preliminary dwarf galaxy detection sensitivity estimates in Drlica-Wagner et al. (2019) indicate that Rubin is sensitive to satellites with $M_{V}<0~{}\mathrm{mag}$ and $\mu<32~{}\mathrm{mag\ arcsec}^{-2}$ in the Southern hemisphere, out to the MW virial radius. However, these projections may be optimistic due to the difficulties of star–galaxy separation; realistic mock simulations of dwarf recovery in Rubin data will be needed to accurately derive its sensitivity. Meanwhile, for M31, Doliva-Dolinsky et al. (2022) show that current PAndAS data is sensitive to systems with $M_{V}\lesssim-5.5~{}\mathrm{mag}$ and $\mu\lesssim 30~{}\mathrm{mag\ arcsec}^{-2}$; for context, Mutlu-Pakdil et al. (2021) find that Rubin is sensitive to even fainter dwarf galaxies at comparable distances, although M31 is not within its field of view.
@@
An assumption underlying our forecasts is that properties of the MW-mass centrals that host the dwarf satellites we consider will be well characterized. Accurate and precise measurements of host halo masses are important for determining expected subhalo and satellite populations, and measurements of secondary properties like host concentration are also helpful (e.g., Mao et al. 2015; Fielder et al. 2019). At present, even the MW’s halo mass remains difficult to measure precisely, although proper motion data from Gaia have significantly reduced this uncertainty (e.g., see Gardner et al. 2021 for a review). For more distant systems, velocity dispersion measurements of globular clusters and of dwarf satellites themselves will likely provide the most precise host mass estimates. Jointly inferring host halo properties and the galaxy–halo connection for their satellites will be necessary to claim a detection of the astrophysical and DM signatures we have explored. Extending our model to marginalize over host halo properties is a natural avenue for future work. We do not expect marginalizing over host properties to change our qualitative conclusions (e.g., the value of combining hosts with different bright satellite populations to break galaxy–halo connection degeneracies); however, our quantitative results concerning the number of hosts needed to achieve given galaxy formation cutoff, WDM, or SHMF constraints will depend on these uncertainties in detail.
We note that our forecasts adopt broad priors on the galaxy–halo connection that are likely to be improved using measurements of brighter dwarf galaxies. As noted above, adopting tighter priors on, e.g., the luminosity function slope and galaxy–halo connection scatter shrinks our systematic error budget and therefore directly improves the precision of our galaxy formation cutoff, WDM, and SHMF constraints. Current surveys like ELVES (Carlsten et al., 2022) and SAGA (Geha et al., 2017; Mao et al., 2021) have already detected large numbers of dwarf satellites throughout the Local Volume and out to $\approx 40~{}\mathrm{Mpc}$, and DESI is supplementing these efforts over a very large area (Darragh-Ford et al., 2023). Combining our framework with data from these surveys will improve galaxy–halo connection constraints (e.g., see Danieli et al. 2023). Furthermore, upcoming facilities including Roman will likely detect large numbers of dwarfs in both the (partially) resolved regimes and in integrated light with distances measured via, e.g., surface brightness fluctuations (Greco et al., 2021). In parallel, ongoing neutral hydrogen surveys like WALLABY are expected to reach HI detection thresholds of $\approx 10^{5}~{}M_{\mathrm{\odot}}$ (Westmeier et al., 2022), probing the gas content of low-mass dwarfs and providing a complementary handle on the faint-end galaxy–halo connection.
@@ 7.2 Long-Term Observational Outlook
In the longer term, next-generation optical/infrared space telescopes have the potential to revolutionize our ability to detect faint dwarf galaxies. For example, The LUVOIR Team (2019) estimates that a $15~{}\mathrm{m}$ optical/infrared space telescope can resolve individual stars out to tens of $\mathrm{Mpc}$ and potentially detect dwarf galaxies with $M_{V}<0~{}\mathrm{mag}$ at distances of $\sim 15~{}\mathrm{Mpc}$. Because of the relatively small field-of-view of such telescope concepts, targeting satellites around known hosts at these distances will be crucial to maximize dwarf galaxy discovery. Such detection sensitivity easily satisfies the assumptions of our forecasts; for example, there are $\approx 30$ MW-mass systems to survey within $12~{}\mathrm{Mpc}$ (e.g., Carlsten et al. 2022). As the community develops the detailed science requirements for the Astro2020-prioritized Habitable Worlds Observatory, the imaging and spectroscopic detection and characterization of faint dwarf galaxies, through both diffuse light and direct stellar-cluster or stellar identifications, is a capability that should be explored. This research will be a key step toward the galaxy formation and DM physics measurements we have forecasted.
@@ 7.3 Modeling Work
Further modeling work will improve our ability to extract galaxy formation and DM physics from future dwarf galaxy surveys. Most importantly, including a broader range of host masses will yield mock dwarf galaxy populations that better span those probed by future surveys. For example, Symphony’s LMC-mass suite contains hosts with $M_{\mathrm{host}}\approx 10^{11}~{}M_{\mathrm{\odot}}$, and its Group suite contains hosts with $M_{\mathrm{host}}\approx 10^{13}~{}M_{\mathrm{\odot}}$. Incorporating these suites will require re-calibrating aspects of our galaxy–halo connection framework that are currently tuned to MW-mass systems, including our subhalo disruption model.
@@
    Even in a dark matter-only setting, the physical origin of correlations between secondary host properties and SHMF amplitude remains unclear. Nadler et al. (2023) show that the mass of the largest subhalo correlates strongly with SHMF amplitude in simulations; there is a similar trend in SAGA data, which shows that the luminosity of the brightest observed satellite correlates with satellite luminosity function amplitude more strongly than central galaxy properties (Mao et al., 2021). For the MW, this correlation partially arises because the LMC brings its own satellites into the MW (e.g., Kallivayalil et al. 2018; Patel et al. 2020). In fact, Nadler et al. (2020b) show that it is difficult to match the spatial anisotropy in the observed MW satellite population without modeling the LMC system. Because the LMC accreted recently and is near pericenter today, this effect is likely transient (e.g., see D’Souza & Bell 2021; Barry et al. 2023). Understanding the origin of the correlation between the mass of the largest subhalo and SHMF amplitude (or the luminosity of the brightest satellite and satellite luminosity function amplitude) would therefore be fruitful. This is particularly relevant because, as shown in Section 4, combining hosts with differing bright satellite abundances can markedly improve galaxy formation constraints.
@@
    Beyond dark matter–only predictions, the presence of a central galaxy has been claimed to affect surviving subhalo and satellite abundances at the factor of $\approx 2$ level (e.g., Garrison-Kimmel et al. 2017b; Richings et al. 2020; Kelley et al. 2019). The strength of this effect is debated (e.g., Webb & Bovy 2020; Green et al. 2022) and may be related to artificial disruption in cosmological simulations more generally (e.g., van den Bosch et al. 2018; van den Bosch & Ogiya 2018; Mansfield et al. 2023). Current hydrodynamic simulations indicate that subhalo disruption is not highly sensitive to central galaxy properties beyond total mass. Because satellite populations will be measured around a variety of central galaxies in the near future, improved theoretical predictions for subhalo and satellite disruption around central galaxies with a range of properties will be helpful. Combining isolated and satellite dwarf populations will also reduce uncertainties associated with subhalo and satellite disruption, which are conservatively marginalized over in our forecasts.
Although our simulations capture a representative range of MW-mass host halos and formation histories (Mao et al., 2015; Nadler et al., 2023), we have not explicitly modeled the impact of host mass and cosmic environment on the galaxy–halo connection. In particular, our galaxy–halo connection model populates subhalos according to the same procedure in all MW-mass hosts we consider. This assumption likely does not capture the complexity of real dwarf galaxy populations, in which galaxy formation and evolution are environmentally dependent (e.g., Christensen et al. 2023; also see Danieli et al. 2023 for discussion in the context of Local Volume satellite populations), particularly for low-mass galaxies affected by photoionization (e.g., Benson et al. 2003). Modeling environmental effects is particularly relevant because the large-scale environment of the Local Volume may be unusual (e.g., Neuzil et al. 2020; McAlpine et al. 2022), thus impacting the expected number of nearby dwarf galaxies and the effects of photoionization on these systems (e.g., Benson et al. 2002; Busha et al. 2010; Li et al. 2014; Dixon et al. 2018). We plan to incorporate host halo property and environmental dependence in our galaxy–halo connection framework in future work.
@@ 8 Prospects for Detecting “Dark” Halos
This peak mass threshold of $3\times 10^{7}~{}M_{\mathrm{\odot}}$ provides a benchmark for the detection of dark (sub)halos using gravitational DM probes. For example, Gilman et al. (2019) predicts that subhalos with present-day masses down to $\approx 10^{7}~{}M_{\mathrm{\odot}}$ can be probed using flux ratio statistics for a sample of $50$ strong gravitational lenses. Facilities including Rubin and Roman are expected to identify thousands of lensed systems (e.g., Oguri & Marshall 2010; Collett 2015); it should be feasible to obtain the deeper follow-up data necessary for DM substructure studies for hundreds of these systems over the next decade, following efforts that are already underway using JWST (Nierenberg et al., 2023). In this context, our forecasts indicate that a statistical detection of dark line-of-sight halos with masses of $\approx 3\times 10^{7}~{}M_{\mathrm{\odot}}$, may be achieved by combining upcoming dwarf galaxy and strong-lensing data. Meanwhile, an unambiguous statistical detection of dark subhalos will likely require sensitivity to substructure with stripped masses below $\approx 10^{7}~{}M_{\mathrm{\odot}}$, although it will still be necessary to model the line-of-sight contribution when interpreting lensing substructure (e.g., Despali et al. 2018; Sengül et al. 2022).
Measuring the gravitational effects of low-mass (sub)halos at an individual level is more challenging. Current gravitational imaging analyses have identified halos with masses of $\approx 10^{9}~{}M_{\mathrm{\odot}}$ (Vegetti et al., 2010, 2012; Hezaveh et al., 2016); due to its excellent angular resolution, very long baseline interferometric (VLBI) observations using the MIT-Green Bank Very Large Array (MG-VLA) are beginning to probe even lower-mass (sub)halos (e.g., Powell et al. 2022). It is therefore plausible that individual (sub)halos below the galaxy formation threshold will be detected in future VLBI data, e.g. from the Next-Generation Very Large Array (ngVLA; Selina et al. 2018; Kadler et al. 2023). Intriguingly, recent JWST observations indicate that highly-magnified and strongly-lensed stars can also potentially probe extremely low-mass DM substructure (Diego et al., 2023).
Low-mass subhalos can also be detected via their effects on stellar streams. Current analyses suggest that the GD-1 stream was impacted by an object with mass below $\approx 10^{8}~{}M_{\mathrm{\odot}}$ (Bonaca et al., 2019) and that power spectra of stream density fluctuations are sensitive to subhalo populations of similar masses (e.g., Banik et al. 2021a). Rubin is expected to increase the precision of such measurements, potentially reaching a perturber mass sensitivity of $\approx 10^{6}~{}M_{\mathrm{\odot}}$ (Drlica-Wagner et al., 2019), with sufficient radial velocity precision. Upcoming stream measurements should provide an independent constraint on the abundance and spatial distribution of low-mass MW subhalos within the next decade.

### v2

@@ 6.3 Interpretation
Our results indicate that future dwarf galaxy surveys can set upper limits on the SHMF at peak subhalo masses from $10^{9}$ to $10^{10}~{}M_{\mathrm{\odot}}$, while simultaneously placing a robust lower limit at $10^{8}~{}M_{\mathrm{\odot}}$. The upper limits on the SHMF amplitude at high masses result from the fact that, if high-mass subhalos are more abundant than in the input CDM model, observed luminosity functions are overpredicted. In particular, it is difficult to “hide” the galaxies that occupy high-mass subhalos by varying galaxy–halo connection parameters without drastically reducing the abundance of lower-mass galaxies and resulting in worse agreement with the mock CDM data (for details, see our analysis of the full SHMF forecast posteriors in Appendix C.3). At lower peak halo masses, and particularly near $M_{\mathrm{peak}}\approx 10^{8}~{}M_{\mathrm{\odot}}$ where our fiducial galaxy formation cutoffs take effect, models with suppressed low-mass subhalo abundances underpredict faint-end luminosity functions; however, models with enhanced low-mass subhalo abundances can be “hidden” by the galaxy formation cutoff. More precise constraints on the mass scale and shape of the cutoff would therefore improve our projected SHMF enhancement measurements.
@@
Our SHMF forecasts can be translated to any DM or early-Universe physics scenario that suppresses or enhances the SHMF relative to CDM. This is most often realized by suppressing or enhancing the formation of (sub)halos relative to CDM; it will therefore be interesting to generalize our modeling framework to directly constrain the linear matter power spectrum, which often dictates the resulting suppression or enhancement of subhalo abundances in such scenarios (e.g., Gilman et al. 2022). Meanwhile, if the SHMF is altered at late times (e.g., via subhalo disruption due to new DM physics), our forecasts are applicable if the abundance of satellite galaxies is correspondingly altered. The connection between subhalo and satellite disruption is nontrivial and requires detailed study in the context of specific DM models (e.g., see Nadler et al. 2020a, 2021a and Mau et al. 2022 for examples in the context of self-interacting and decaying DM); we leave studies of such scenarios to future work.
@@
Our SHMF suppression results are relevant for a variety of warm, interacting, and ultralight DM models (e.g., Bechtol et al. 2022). Meanwhile, DM models that enhance the SHMF more easily evade constraints from dwarf galaxy population statistics because their effects can be counteracted by galaxy formation physics. Such models include, e.g., ultralight axions that undergo parametric resonance in the early Universe (Arvanitaki et al., 2020; Cyncynates et al., 2022) or DM produced during a period of early matter domination (Erickcek & Sigurdson, 2011). We expect that dwarf galaxy observables beyond luminosity, size, and distance (e.g., velocity dispersions; Kim & Peter 2021; Esteban et al. 2023) will help probe DM physics that enhances (sub)halo abundances. Distinct small-scale structure probes like strong gravitational lensing and stellar streams, which are more sensitive to subhalos density profiles than dwarf galaxy population statistics, will complement this effort.
@@ 7.1 Near-term Observational Capabilities
Our assumption that complete satellite populations of several MW-mass systems can be detected is clearly optimistic given current and near-future observational capabilities. In particular, this level of completeness will be difficult to achieve even for the MW and M31 in the next decade. Preliminary dwarf galaxy detection sensitivity estimates in Drlica-Wagner et al. (2019) indicate that Rubin is sensitive to satellites with $M_{V}<0~{}\mathrm{mag}$ and $\mu<32~{}\mathrm{mag\ arcsec}^{-2}$ in the Southern Hemisphere, out to the MW virial radius. However, these projections may be optimistic due to the difficulties of star–galaxy separation; realistic mock simulations of dwarf recovery in Rubin data will be needed to accurately derive its sensitivity. Meanwhile, for M31, Doliva-Dolinsky et al. (2022) showed that current PAndAS data is sensitive to systems with $M_{V}\lesssim-5.5~{}\mathrm{mag}$ and $\mu\lesssim 30~{}\mathrm{mag\ arcsec}^{-2}$; for context, Mutlu-Pakdil et al. (2021) found that Rubin is sensitive to even fainter dwarf galaxies at comparable distances, although M31 is not within its field of view.
@@
An assumption underlying our forecasts is that properties of the MW-mass centrals that host the dwarf satellites we consider will be well characterized. Accurate and precise measurements of host halo masses are important for determining expected subhalo and satellite populations, and measurements of secondary properties like host concentration are also helpful (e.g., Mao et al. 2015; Fielder et al. 2019). At present, even the MW’s halo mass remains difficult to measure precisely; although, proper motion data from Gaia have significantly reduced this uncertainty (e.g., see Gardner et al. 2021 for a review). For more distant systems, velocity dispersion measurements of globular clusters and of dwarf satellites themselves will likely provide the most precise host mass estimates. Jointly inferring host halo properties and the galaxy–halo connection for their satellites will be necessary to claim a detection of the astrophysical and DM signatures we have explored. Extending our model to marginalize over host halo properties is a natural avenue for future work. We do not expect marginalizing over host properties to change our qualitative conclusions (e.g., the value of combining hosts with different bright satellite populations to break galaxy–halo connection degeneracies); however, our quantitative results concerning the number of hosts needed to achieve given galaxy formation cutoff, WDM, or SHMF constraints will depend on these uncertainties in detail.
We note that our forecasts adopt broad priors on the galaxy–halo connection that are likely to be improved using measurements of brighter dwarf galaxies. As noted above, adopting tighter priors on, e.g., the luminosity function slope and galaxy–halo connection scatter shrinks our systematic error budget and therefore directly improves the precision of our galaxy formation cutoff, WDM, and SHMF constraints. Current surveys like ELVES (Carlsten et al., 2022) and SAGA (Geha et al., 2017; Mao et al., 2021) have already detected large numbers of dwarf satellites throughout the Local Volume and out to $\approx 40~{}\mathrm{Mpc}$, and DESI is supplementing these efforts over a very large area (Darragh-Ford et al., 2023). Combining our framework with data from these surveys will improve galaxy–halo connection constraints (e.g., Danieli et al. 2023). Furthermore, upcoming facilities including Roman will likely detect large numbers of dwarfs in both the (partially) resolved regimes and in integrated light with distances measured via, e.g., surface brightness fluctuations (Greco et al., 2021). In parallel, ongoing neutral hydrogen surveys like WALLABY are expected to reach HI detection thresholds of $\approx 10^{5}~{}M_{\mathrm{\odot}}$ (Westmeier et al., 2022), probing the gas content of low-mass dwarfs and providing a complementary handle on the faint-end galaxy–halo connection.
@@ 7.2 Long-term Observational Outlook
In the longer term, next-generation optical/infrared space telescopes have the potential to revolutionize our ability to detect faint dwarf galaxies. For example, The LUVOIR Team (2019) estimated that a $15~{}\mathrm{m}$ optical/infrared space telescope can resolve individual stars out to tens of megaparsecs and potentially detect dwarf galaxies with $M_{V}<0~{}\mathrm{mag}$ at distances of $\sim 15~{}\mathrm{Mpc}$. Because of the relatively small field of view of such telescope concepts, targeting satellites around known hosts at these distances will be crucial to maximize dwarf galaxy discovery. Such detection sensitivity easily satisfies the assumptions of our forecasts; for example, there are $\approx 30$ MW-mass systems to survey within $12~{}\mathrm{Mpc}$ (e.g., Carlsten et al. 2022). As the community develops the detailed science requirements for the Astro2020-prioritized Habitable Worlds Observatory, the imaging and spectroscopic detection and characterization of faint dwarf galaxies, through both diffuse light and direct stellar-cluster or stellar identifications, is a capability that should be explored. This research will be a key step toward the galaxy formation and DM physics measurements we have forecasted.
@@ 7.3 Modeling Work
Further modeling work will improve our ability to extract galaxy formation and DM physics from future dwarf galaxy surveys. Most importantly, including a broader range of host masses will yield mock dwarf galaxy populations that better span those probed by future surveys. For example, Symphony’s LMC-mass suite contains hosts with $M_{\mathrm{host}}\approx 10^{11}~{}M_{\mathrm{\odot}}$, and its Group suite contains hosts with $M_{\mathrm{host}}\approx 10^{13}~{}M_{\mathrm{\odot}}$. Incorporating these suites will require recalibrating aspects of our galaxy–halo connection framework that are currently tuned to MW-mass systems, including our subhalo disruption model.
@@
    Even in a dark-matter-only setting, the physical origin of correlations between secondary host properties and SHMF amplitude remains unclear. Nadler et al. (2023) showed that the mass of the largest subhalo correlates strongly with SHMF amplitude in simulations; there is a similar trend in SAGA data, which shows that the luminosity of the brightest observed satellite correlates with satellite luminosity function amplitude more strongly than central galaxy properties (Mao et al., 2021). For the MW, this correlation partially arises because the LMC brings its own satellites into the MW (e.g., Kallivayalil et al. 2018; Patel et al. 2020). In fact, Nadler et al. (2020b) showed that it is difficult to match the spatial anisotropy in the observed MW satellite population without modeling the LMC system. Because the LMC accreted recently and is near pericenter today, this effect is likely transient (e.g., D’Souza & Bell 2021; Barry et al. 2023). Understanding the origin of the correlation between the mass of the largest subhalo and SHMF amplitude (or the luminosity of the brightest satellite and satellite luminosity function amplitude) would therefore be fruitful. This is particularly relevant because, as shown in Section 4, combining hosts with differing bright satellite abundances can markedly improve galaxy formation constraints.
@@
    Beyond dark-matter-only predictions, the presence of a central galaxy has been claimed to affect surviving subhalo and satellite abundances at the factor of $\approx 2$ level (e.g., Garrison-Kimmel et al. 2017b; Kelley et al. 2019; Richings et al. 2020). The strength of this effect is debated (e.g., Webb & Bovy 2020; Green et al. 2022) and may be related to artificial disruption in cosmological simulations more generally (e.g., van den Bosch & Ogiya 2018; van den Bosch et al. 2018; Mansfield et al. 2023). Current hydrodynamic simulations indicate that subhalo disruption is not highly sensitive to central galaxy properties beyond total mass. Because satellite populations will be measured around a variety of central galaxies in the near future, improved theoretical predictions for subhalo and satellite disruption around central galaxies with a range of properties will be helpful. Combining isolated and satellite dwarf populations will also reduce uncertainties associated with subhalo and satellite disruption, which are conservatively marginalized over in our forecasts.
Although our simulations capture a representative range of MW-mass host halos and formation histories (Mao et al., 2015; Nadler et al., 2023), we have not explicitly modeled the impact of host mass and cosmic environment on the galaxy–halo connection itself. In particular, our galaxy–halo connection model populates subhalos according to the same procedure in all MW-mass hosts we consider. This assumption likely does not capture the complexity of real dwarf galaxy populations, in which galaxy formation and evolution are environmentally dependent (e.g., Christensen et al. 2023; also see Danieli et al. 2023 for discussion in the context of Local Volume satellite populations), particularly for low-mass galaxies affected by photoionization (e.g., Benson et al. 2003). Modeling environmental effects is particularly relevant because the large-scale environment of the Local Volume may be unusual (e.g., Neuzil et al. 2020; McAlpine et al. 2022), thereby impacting the expected number of nearby dwarf galaxies and the effects of photoionization on these systems (e.g., Benson et al. 2002; Busha et al. 2010; Li et al. 2014; Dixon et al. 2018).
We plan to incorporate host halo property and environmental dependence in our galaxy–halo connection framework in future work. In a similar spirit, applying our inference framework to dwarf galaxy populations predicted by hydrodynamic simulations and SAMs, including in beyond-CDM scenarios—rather than mock data generated by applying our empirical galaxy–halo connection model to CDM simulations—would be a useful test of our framework’s flexibility.
@@ 8 Prospects for Detecting “Dark” Halos
This peak mass threshold of $3\times 10^{7}~{}M_{\mathrm{\odot}}$ provides a benchmark for the detection of dark (sub)halos using gravitational DM probes. For example, Gilman et al. (2019) predicted that subhalos with present-day masses down to $\approx 10^{7}~{}M_{\mathrm{\odot}}$ can be probed using flux ratio statistics for a sample of $50$ strong gravitational lenses. Facilities including Rubin and Roman are expected to identify thousands of strong lenses in the coming years (e.g., Oguri & Marshall 2010; Collett 2015); it should be feasible to obtain the deeper follow-up data necessary for DM substructure studies for hundreds of these systems over the next decade, following efforts that are already underway using JWST (Nierenberg et al., 2023). In this context, our forecasts indicate that a statistical detection of dark line-of-sight halos with masses of $\approx 3\times 10^{7}~{}M_{\mathrm{\odot}}$, may be achieved by combining upcoming dwarf galaxy and strong-lensing data. Meanwhile, an unambiguous statistical detection of dark subhalos will likely require sensitivity to substructure with stripped masses below $\approx 10^{7}~{}M_{\mathrm{\odot}}$, although it will still be necessary to model the line-of-sight contribution when interpreting lensing substructure (e.g., Despali et al. 2018; Sengül et al. 2022).
Measuring the gravitational effects of low-mass (sub)halos at an individual level is more challenging. Current gravitational imaging analyses have identified halos with masses of $\approx 10^{9}~{}M_{\mathrm{\odot}}$ (Vegetti et al., 2010, 2012; Hezaveh et al., 2016); due to its excellent angular resolution, very long baseline interferometric (VLBI) observations using the MIT-Green Bank Very Large Array are beginning to probe even lower-mass (sub)halos (e.g., Powell et al. 2022). It is therefore plausible that individual (sub)halos below the galaxy formation threshold will be detected in future VLBI data, e.g. from the Next-Generation Very Large Array (Selina et al., 2018; Kadler et al., 2023). Intriguingly, recent JWST observations indicate that highly magnified and strongly lensed stars can also potentially probe extremely low-mass DM substructure (Diego et al., 2023).
Low-mass subhalos can also be detected via their gravitational effects on stellar streams. For example, current analyses suggest that the GD-1 stream was impacted by an object with mass below $\approx 10^{8}~{}M_{\mathrm{\odot}}$ (Bonaca et al., 2019) and that power spectra of stream density fluctuations are sensitive to subhalo populations of similar masses (e.g., Banik et al. 2021a). Rubin is expected to increase the precision of such measurements, potentially reaching a perturber mass sensitivity of $\approx 10^{6}~{}M_{\mathrm{\odot}}$ (Drlica-Wagner et al., 2019). Upcoming stream measurements should provide an independent constraint on the abundance and spatial distribution of low-mass MW subhalos within the next decade.